OpenStudy (anonymous):

Choose the correct product of (5x − 3)(5x + 3). 25x2 + 30x + 9 my answer

OpenStudy (mathstudent55):

Do you know FOIL?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Can you do it and show every step without combining like terms?

OpenStudy (whpalmer4):

if the last two numbers have opposite signs, your answer must also have a negative sign for the last term... \[(a+b)(a-b) = a^2+ab-ab-b^2\]

OpenStudy (whpalmer4):

"last two numbers" being the constant terms in the product, namely "-3" and "+3"

OpenStudy (anonymous):

\[\left( 5x+3\left( 5x-3 \right) \right)=\left( 5x ^{2} \right)-\left( 3 \right)^{2}=25x ^{2 }-9\]

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

\[\left( a+b \right)\left( a-b \right)=a ^{2}-b ^{2}\]

OpenStudy (whpalmer4):

Learn to recognize that pattern, called a difference of squares. Knowing it, you can instantly write the answer to this problem: \[(5x-3)(5x+3) = (5x)^2 -(3)^2 = 25x^2-9\] and you can also instantly factor \[25x^2-9 = (5x-3)(5x+3)\]